OpenStudy (anonymous):

Solve for x: 3 over x minus 2 over 3 equals 5 over 6

OpenStudy (anonymous):

isi it \(\frac{ 3 }{ \frac{ x-2 }{ 3 } } = \frac{ 5 }{ 6 }\)..?

OpenStudy (anonymous):